OpenStudy (anonymous):

Find the line that is tangent to the circle x^2+y^2=25 at the point 3,-4.

OpenStudy (mathmate):

The centre is at (0,0), the radial line passing through the centre (origin) has a slope of (-4/3), so the slope of the tangent is 3/4. A line through (3,-4) with slope 3/4 is therefore y=(3/4)(x-3)+(-4)

OpenStudy (amistre64):

since 3,-4 is the upper part of a circle; sqrt(yada)

OpenStudy (amistre64):

3 over 4 down, doh!! under side

OpenStudy (amistre64):

-sqrt(yada)

OpenStudy (anonymous):

and how do we know what the tangent slope is? start with \[x^2+y^2=25\] then \[2x+2yy'=0\] \[y'=-\frac{x}{y}\] and so at \[(3,-4)\] the slope is \[-\frac{3}{-4}=\frac{3}{4}\]

OpenStudy (anonymous):

now point slope formula gives it

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

that was implicit diff that i used. you can solve for y if you like and then take the derivative there. you will get \[y'=\pm\frac{x}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

but I'm in the parametric equations of line section? we haven't been using any derivatives

OpenStudy (mathmate):

You don't need any derivative to find the slope of the radial line and the tangent line. You only need the relationship m1*m2 = -1. ] Follow the reasoning of my very first response.